11. A small accounting firm pays each of its five clerks S50,000, two junior accountants S100,000, and one senior accountant $150,000. (a) What is the mean salary? (b) Suppose all eight employees receive a $10,000 raise. What is the mean salary? 12. Suppose a set of temperatures measured in celsius has a mean 30 C and standard deviation 5°C. What is the mean and standard deviation of the same set of temperatures measured in fahrenheit? The relationship between fahrenheit (F) and celsius (C) is FC+32
Answer #1

Solution: 11. A small accounting firm pays each of its five clerks $50,000, two junior accountants $100,000, and one senior accountant $150,000

(a) What is the mean salary?

Answer: We are given that there are five clerks, two junior accountants and one senior accountant. Therefore there are 8 employees in a small accounting firm.

The total amount that a small accounting firm pays to its eight employees is:

5 × 50000+2 × 100000+ 1 × 150000-250000+200000+ 150000 600000

Mean salary is

\frac{600,000}{8}=75,000

Therefore the mean salary is \ 75,000

(b) Suppose all eight employees receive a $10,000 raise. What is the mean salary?

Answer: If all eight employees receive a $10,000 raise, then their salary becomes.

Salary of 5 clerks =5 \times (50000+10000)=5 \times 60000=300000

Salary of 2 junior accountants =2 \times (100000+10000)=2 \times 110000=220000

Salary of senior accountant =1 \times (150000+10000)=160000

Now the mean salary is:

\frac{300000+220000+160000}{8}=85000

Therefore the mean salary after the all eight employees receive a $10,000 raise is $85,000.

12. Suppose a set of temperatures measured in celsius has a mean 30^{\cdot } C and standard deviation 5^{\cdot } C. What is the mean and standard deviation of the same set of temperatures measured in fahrenheit? The relationship between fahrenheit (F) and celsius (C) is

F=\left ( \frac{9}{5} \right )C+32

Answer: The mean in fahrenheit is:

Mean=\left ( \frac{9}{5} \right )30+32

=54+32

=86

Therefore the mean temperature in fahrenheit is 86^{\cdot }F

The standard deviation in fahrenheit is:

Standard-deviation=\left ( \frac{9}{5} \right )5+32

=9+32

=41

Therefore the standard deviation in fahrenheit is 41^{\cdot }F